OpenStudy (harrisonshark):

x2 + 8x - 5 = 0 Solve this quadratic equation using the quadratic formula.

OpenStudy (anonymous):

pls state the quadratic formula below, thnx

OpenStudy (harrisonshark):

x2 + 8x - 5 = 0

OpenStudy (anonymous):

"Solve this quadratic equation using the quadratic formula." "x2 + 8x - 5 = 0" is the "quadratic equation"

OpenStudy (aric200):

X=1 Third

OpenStudy (harrisonshark):

thank you

OpenStudy (harrisonshark):

kx your a d

OpenStudy (anonymous):

\[\huge x = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] This is the quadratic formula.

OpenStudy (anonymous):

@aric200 let's teach him how to get the answer first, pls :)

OpenStudy (anonymous):

It's just plug and chug.

OpenStudy (aric200):

<3 Sorry :P

OpenStudy (harrisonshark):

kx lets not be a douche first

OpenStudy (anonymous):

thanks @harrisonshark u r very kind ;)

OpenStudy (aric200):

No offence but should you really be giving help on here when you are spelling you u and are as r?

OpenStudy (anonymous):

He's a douche for saying not just give the answer but explain how to get there? Ok.

OpenStudy (anonymous):

Well I already know this is going no where, so bye.

OpenStudy (harrisonshark):

nono @iambatman he cprrected me at the start6

OpenStudy (harrisonshark):

wait batmans wrong

OpenStudy (anonymous):

The standard form of a quadratic equation looks like: ax^2+bx+c = 0, you know a,b, and c. And I showed you the formula earlier. So give it a try.

OpenStudy (harrisonshark):

it's not because there is no letters

OpenStudy (anonymous):